# Digital Linearizer Based on 1-Bit Quantizations

### v1

@@ Abstract
Index Terms:
Analog-to-digital interfaces, nonlinear distortion, memoryless linearizer, 1-bit quantization.
@@ II Distortion and Linearization
In total, this linearizer requires $12K-$ 1 multiplications and $K$ additions per corrected output sample, with $K$ multiplications to produce $d_{k}v^{k}(n)$ and $K-1$ multiplications for all the $v^{k}(n)$ computations. Alternatively, one may use the so-called Wiener linearizer [5] where the multiplications by $d_{k}$, $k=2,3,\ldots,K$, seen in Fig. 1, are carried out before the operations $v^{k}(n)$, but it does not change the implementation complexity and performance.
@@
with $f_{m}$, $m=1,2,\ldots,N$, representing a set of $N$ nonlinear functions. In [2], the ReLU operations $f_{m}(v(n))=\max\{0,v(n)+b_{m}\}$ or modulus operations $f_{m}(v(n)+b_{m})=|v(n)+b_{m}|$ were used due to their simplicity in hardware implementations [6]. Further, the bias values $b_{m}$ were uniformly distributed between $-b_{\max}$ and $b_{\max}$, i.e., $b_{m}=-b_{\max}+2(m-1)b_{\max}/(N-1)$, where the value of $b_{\max}$ was determined in the design. It was demonstrated in [2], that the linearizer in Fig. 2 can outperform the Hammerstein linearizer even when the nonlinearities have been generated through the memoryless polynomial model in (1). However, the linearizer in Fig. 2 still requires $N+1$ multiplications and $12N+$ 1 additions per corrected output sample.
@@ III Proposed Linearizer
The bias values $b_{m}$ are then again uniformly distributed between $-b_{\max}$ and $b_{\max}$. Utilizing (5), $b_{m}$ are thus chosen as
@@
The implication of the choices in (4)–(6) is two-fold. Firstly, for each sample index $n$, since the output of each nonlinear operation $f_{m}\big{(}v(n)+b_{m}\big{)}$ is either one or zero, and $b_{m}$ are fixed and uniformly distributed between $-b_{\max}$ and $b_{\max}$, there are $N+1$ possible combinations of zeros and ones in the $N$ inputs to $w_{m}$, and these combinations contain consecutive zeros followed by consecutive ones (except for two combinations with only zeros or only ones). For example, with $N=4$, the only possible combinations are $0000$, $0001$, $0011$, $0111$, and $1111$. This means that the $N$ multiplications by $w_{m}$ followed by the corresponding $N-1$ additions, and the addition of $c_{0}$, can be implemented with a look-up table of size $N+1$ and with entries $u_{q}$, $q=0,1,\ldots,N$, where
@@
Secondly, with $b_{\max}$ as in (5), and with $b_{m}$ uniformly distributed between $-b_{\max}$ and $b_{\max}$, it is readily shown that the whole signal-value region from $-1$ to $1$, of the input signal $v(n)$, can be divided into $N+1$ consecutive sub-regions of the same size $2/(N+1)$, with each sub-region corresponding to a unique combination of the $N+1$ possible zero/one combinations at the quantizers’ outputs. This is exemplified in Fig. 3 for $N=7$, and it means in general that neither the $N$ additions by the bias values $b_{m}$ nor the $1$-bit quantizations and subsequent multiplications have to be carried out. Instead, the look-up table can be directly addressed by identifying and mapping the input signal level to a corresponding memory address. For this, it suffices to use $\lceil\log_{2}(N+1)\rceil$ memory-address bits as exemplified in Fig. 3. The proposed linearizer can therefore be implemented as Fig. 4 depicts [^2] By scaling the linear and nonlinear branches by $1/c_{1}$, one could remove the multiplication in the linear branch. However, in a practical implementation, scaling of the signal levels is required [7], which means that one multiplication is generally needed anyhow.. It can be equivalently described by the scheme in Fig. 5, since the look-up table based implementation corresponds to adding a time-varying value $w\big{(}v(n)\big{)}$ which can take on the $N+1$ values in the table and whose value for each $n$ is determined by the value of $v(n)$. Before proceeding, it is noted that post-correction methods for ADCs utilizing look-up tables have appeared earlier, as in [8] for the memoryless case. However, those methods are different as they target small-scale nonlinearities caused by nonideal quantizers (deviations from uniform quantizations) and the correction table is then designed based on ADC output-code statistics. Further, those correction methods (with one fixed table) typically offer modest signal-to-noise-and-distortion ratio (SNDR) improvements for some frequencies but can even deteriorate the SNDR for other frequencies [8]. The linearizer in this paper targets more general nonlinearities and larger SNDR improvements, and it is designed to correct all frequencies in a prespecified frequency band. To further stress the difference, it is noted that 1) the proposed linearizer is designed via the basic structure in Fig. 2 and it can also be implemented in that way (but the look-up table implementation can be more efficient), which is not the case for the previous look-up-table-based methods, and 2) the size of the memory ($N+1$) in the proposal can be kept relatively small as it is determined by the number of nonlinear branches ($N$), and not by the number of data-bits combinations utilized when designing the tables in the previous methods.
@@ IV Evaluations and Comparisons
in which case the signal corresponds to the quadrature (imaginary) part of $31$ active subcarriers in a $64$-subcarrier OFDM signal with a random frequency offset $\Delta\omega$. In the design and evaluation we use, respectively, $R=1$ and $M=2500$ signals with randomly generated frequency offsets assuming uniform distribution between $-\pi/64$ and $\pi/64$, quantized to $8$ bits, and of length $L=8192$. The gain $G$ is selected so that the distorted signal is below one in magnitude. For the $L_{2}$-regularization, we use $\lambda=0.0002$. Figure 6 plots the spectrum before and after linearization for one of the signals. Figure 7 plots the mean SNDR over 2500 signals for each linearizer instance (with an SNDR variance of some 0.5 dB for all instances) versus the number of branches for the proposed linearizer, the one in [2], and the Hammerstein linearizer, designed in the same least-squares sense (for all designs, the optimized parameter values were quantized to $12$ bits). For the signals considered in this example, quantized to $8$ bits, the signal-to-noise ratio (SNR) is some $42$ dB without distortion, and the SNDR is some $25$ dB for the distorted signal before the linearization. Hence, the linearizer can at most improve the SNDR by $17$ dB in this example, which corresponds to almost $3$ bits improvement.
As seen in Fig. 7, for the proposed linearizer, the SNDR does not increase monotonically. This is because its bias values are not optimized but instead selected according to (5) and (6) to enable the look-up table implementation. Further, the proposed linearizer requires more nonlinear branches than the other two methods to reach the same SNDR. However, to assess the implementation complexity, one should not consider the number of branches but instead the complexity required to linearize each output sample. With $N$ nonlinear branches, the Hammerstein linearizer (where $N=K-1)$ requires $12N+$ 1 multiplications and $N+1$ additions whereas the linearizer in [2] requires $N+1$ multiplications and $12N+$ 1 additions. In general, multiplications are considerably more complex to implement than additions and the linearizer in [2] is thus more efficient than the Hammerstein linearizer for the same number of nonlinear branches. The proposed linearizer, implemented with a look-up table, requires only one multiplication and one addition per output sample. Hence, in terms of computational complexity, this linearizer clearly outperforms the other two, as illustrated in Fig. 8 which plots the number of multiplications versus the SNDR for the three methods. However, the proposed linearizer also have the additional cost of the memory look-up implementation, which is dependent upon the hardware platform.

### v2

@@ II Distortion and Linearization
In total, this linearizer requires 2K12K-1 multiplications and $K$ additions per corrected output sample, with $K$ multiplications to produce $d_{k}v^{k}(n)$ and K1K-1 multiplications for all the $v^{k}(n)$ computations. Alternatively, one may use the so-called Wiener linearizer [5] where the multiplications by $d_{k}$, $k=2,3,\ldots,K$, seen in Fig. 1, are carried out before the operations $v^{k}(n)$, but it does not change the implementation complexity and performance.
@@
with $f_{m}$, $m=1,2,\ldots,N$, representing a set of $N$ nonlinear functions. In [2], the ReLU operations $f_{m}(v(n))=\max\{0,v(n)+b_{m}\}$ or modulus operations $f_{m}(v(n)+b_{m})=|v(n)+b_{m}|$ were used due to their simplicity in hardware implementations [6]. Further, the bias values $b_{m}$ were uniformly distributed between $b_{\max}$ and $b_{\max}$, i.e., $b_{m}=-b_{\max}+2(m-1)b_{\max}/(N-1)$, where the value of $b_{\max}$ was determined in the design. It was demonstrated in [2], that the linearizer in Fig. 2 can outperform the Hammerstein linearizer even when the nonlinearities have been generated through the memoryless polynomial model in (1). However, the linearizer in Fig. 2 still requires $N+1$ multiplications and $2N+1$ additions per corrected output sample.
@@ III Proposed Linearizer
The bias values $b_{m}$ are then again uniformly distributed between $b_{\max}$ and $b_{\max}$. Utilizing (5), $b_{m}$ are thus chosen as
@@
The implication of the choices in (4)–(6) is two-fold. Firstly, for each sample index $n$, since the output of each nonlinear operation $f_{m}\big(v(n)+b_{m}\big)$ is either one or zero, and $b_{m}$ are fixed and uniformly distributed between $b_{\max}$ and $b_{\max}$, there are $N+1$ possible combinations of zeros and ones in the $N$ inputs to $w_{m}$, and these combinations contain consecutive zeros followed by consecutive ones (except for two combinations with only zeros or only ones). For example, with $N=4$, the only possible combinations are $0000$, $0001$, $0011$, $0111$, and $1111$. This means that the $N$ multiplications by $w_{m}$ followed by the corresponding N1N-1 additions, and the addition of $c_{0}$, can be implemented with a look-up table of size $N+1$ and with entries $u_{q}$, $q=0,1,\ldots,N$, where
@@
Secondly, with $b_{\max}$ as in (5), and with $b_{m}$ uniformly distributed between $b_{\max}$ and $b_{\max}$, it is readily shown that the whole signal-value region from 1-1 to $1$, of the input signal $v(n)$, can be divided into $N+1$ consecutive sub-regions of the same size $2/(N+1)$, with each sub-region corresponding to a unique combination of the $N+1$ possible zero/one combinations at the quantizers’ outputs. This is exemplified in Fig. 3 for $N=7$, and it means in general that neither the $N$ additions by the bias values $b_{m}$ nor the $1$-bit quantizations and subsequent multiplications have to be carried out. Instead, the look-up table can be directly addressed by identifying and mapping the input signal level to a corresponding memory address. For this, it suffices to use $\lceil\log_{2}(N+1)\rceil$ memory-address bits as exemplified in Fig. 3. The proposed linearizer can therefore be implemented as Fig. 4 depicts [^2] By scaling the linear and nonlinear branches by $1/c_{1}$, one could remove the multiplication in the linear branch. However, in a practical implementation, scaling of the signal levels is required [7], which means that one multiplication is generally needed anyhow.. It can be equivalently described by the scheme in Fig. 5, since the look-up table based implementation corresponds to adding a time-varying value $w\big(v(n)\big)$ which can take on the $N+1$ values in the table and whose value for each $n$ is determined by the value of $v(n)$. Before proceeding, it is noted that post-correction methods for ADCs utilizing look-up tables have appeared earlier, as in [8] for the memoryless case. However, those methods are different as they target small-scale nonlinearities caused by nonideal quantizers (deviations from uniform quantizations) and the correction table is then designed based on ADC output-code statistics. Further, those correction methods (with one fixed table) typically offer modest signal-to-noise-and-distortion ratio (SNDR) improvements for some frequencies but can even deteriorate the SNDR for other frequencies [8]. The linearizer in this paper targets more general nonlinearities and larger SNDR improvements, and it is designed to correct all frequencies in a prespecified frequency band. To further stress the difference, it is noted that 1) the proposed linearizer is designed via the basic structure in Fig. 2 and it can also be implemented in that way (but the look-up table implementation can be more efficient), which is not the case for the previous look-up-table-based methods, and 2) the size of the memory ($N+1$) in the proposal can be kept relatively small as it is determined by the number of nonlinear branches ($N$), and not by the number of data-bits combinations utilized when designing the tables in the previous methods.
@@ IV Evaluations and Comparisons
in which case the signal corresponds to the quadrature (imaginary) part of $31$ active subcarriers in a $64$-subcarrier OFDM signal with a random frequency offset $\Delta\omega$. In the design and evaluation we use, respectively, $R=1$ and $M=2500$ signals with randomly generated frequency offsets assuming uniform distribution between $\pi/64$ and $\pi/64$, quantized to $8$ bits, and of length $L=8192$. The gain $G$ is selected so that the distorted signal is below one in magnitude. For the $L_{2}$-regularization, we use $\lambda=0.0002$. Figure 6 plots the spectrum before and after linearization for one of the signals. Figure 7 plots the mean SNDR over 2500 signals for each linearizer instance (with an SNDR variance of some 0.5 dB for all instances) versus the number of branches for the proposed linearizer, the one in [2], and the Hammerstein linearizer, designed in the same least-squares sense (for all designs, the optimized parameter values were quantized to $12$ bits). For the signals considered in this example, quantized to $8$ bits, the signal-to-noise ratio (SNR) is some $42$ dB without distortion, and the SNDR is some $25$ dB for the distorted signal before the linearization. Hence, the linearizer can at most improve the SNDR by $17$ dB in this example, which corresponds to almost $3$ bits improvement.
As seen in Fig. 7, for the proposed linearizer, the SNDR does not increase monotonically. This is because its bias values are not optimized but instead selected according to (5) and (6) to enable the look-up table implementation. Further, the proposed linearizer requires more nonlinear branches than the other two methods to reach the same SNDR. However, to assess the implementation complexity, one should not consider the number of branches but instead the complexity required to linearize each output sample. With $N$ nonlinear branches, the Hammerstein linearizer (where N=K1)N=K-1) requires $2N+1$ multiplications and $N+1$ additions whereas the linearizer in [2] requires $N+1$ multiplications and $2N+1$ additions. In general, multiplications are considerably more complex to implement than additions and the linearizer in [2] is thus more efficient than the Hammerstein linearizer for the same number of nonlinear branches. The proposed linearizer, implemented with a look-up table, requires only one multiplication and one addition per output sample. Hence, in terms of computational complexity, this linearizer clearly outperforms the other two, as illustrated in Fig. 8 which plots the number of multiplications versus the SNDR for the three methods. However, the proposed linearizer also have the additional cost of the memory look-up implementation, which is dependent upon the hardware platform.